MOBY-L274 experiment with MOBYrefresh - BS03cfg005 - test scans with BS03cfg005 after its MOBY265 deployment (RS04 is attached) - NEW PROCESSING

BS03 postcal data collected after deployment on MOBY265 with the RS04.

While collecting the data Mike noticed, after 2 days of trying to figure out what was wrong, that Es is Track 13 not 2 and all the other graphs are reverse from what we thought. The RS track numbers are correct. So I have reprocessed the MOBY265 with the correct track numbers and am now fixing this data set.

BLUE jumper cables RED jumper cables
Track Sensors Old wrong number Track Sensors  
13 Es 2 2 Es  
6 LuTop 6 6 LuTop  
9 LuMid 9 9 LuMid  
2 LuBot 13 13 LuBot  

Problems seen (and solved) and good things:

  1. Tracks where not on the sensors we thought when Mike started looking at data. I have corrected the track numbers for the BS here and in MOBY265. You can see the old and new track definitions in the table above.
  2. Looks like the track definitions have moved a bit from the MOBY263 tracks.
  3. Track 6 Lu system response has an odd shape compared to the other Lu trcks.
  4. Wavelength cals where done on Es, LuTop and LuBot
Track pixels:     1       2       3       4       5       6       7       8       9      10      11      12      13      14  
Left:             1      78     151     229     297     367     439     516     581     659     728     800     873     948  
Right:           77     150     228     296     366     438     515     580     658     727     799     872     947    1023  

Log Sheets

MOBY265 postcal Es\Lu rsp

Wavelength cals
P(2,:) = [-3.62437e-09 1.713853e-05 0.3370589 337.6754];
P(6,:) = [-3.642533e-09 1.710183e-05 0.337196 337.5819];
P(13,:) = [-2.567331e-09 1.534803e-05 0.3378738 337.7046];

All the raw data files and their images , Table of KEYWORDS variable
Page Number
Link
Description
Date
1.01
Day 01 - 28 Nov Raw Data
DAY01 - led testing (light is not showing up on the right tracks)
30 Nov 2018
1.02
Day 02 - 29 Nov Raw Data
DAY02 - Es track 2 postcal sysrsp (all are dark)
30 Nov 2018
1.03
Day 03 - 30 Nov Raw Data
DAY03 - Es postcal sysrsp (Es is showing up on Track 13 not 2 and vis versa) other testing - Es rsp synced with RS
30 Nov 2018
1.04
Day 04 - 4 Dec Raw Data
DAY04 - LuTop, Mid and Bot postcal sysrsp - Lu rsp synced with RS
5 Dec 2018
1.05
Day 05 - 5 Dec Raw Data
DAY05 - Es wavecals
5 Dec 2018
1.06
Day 06 - 6 Dec Raw Data
DAY06 - LuTop/Bot wavecals
6 Dec 2018
1.07
Day 07 - 7 Dec Raw Data
DAY07 - LuTop int time cals
10 Dec 2018
Daily data sets
2.01 check track defs Check the track definitions for the Es and Lu tracks 4 Dec 2018
Day 3 data
13.01 Es system response data Es system response data from Track 13 on day 3 using F601(c3) lamp file - need new wavecal to finalize 4 Dec 2018
Day 4 data
14.02 Es/Lu system response Es/Lu system response data from day 3 and 4 - need new wavecal to finalize 5 Dec 2018
Day 5 data
15.02 Es wavlength cal Es wavecals (HgA,Ne) 5 Dec 2018
Day 6 data
16.01 LuTop wavlength cal LuTop wavecals (all lamps) 6 Dec 2018
16.02 LuBot wavlength cal LuBot wavecals (HgA,Ne) 6 Dec 2018
Day 7 data - LuTop int time cal
17.01 int time cals LuTop int time cals (Brite, Mid, dim and dimmer) 10 Dec 2018
17.02 int time cals 10 sec LuTop int time cals (Brite, Mid, dim and dimmer) - normalized to 10 sec 21 May 2019
17.03 compare 1 and 10 sec cal Comparing the 1 and 10 sec normalize int time cal 22 May 2019
17.04 SNR SNR of the int time cal data 29 May 2019
Compilation of day data
20.01 UV nets->BS link RS pow lix data are negative- We found some nevative values in the low pixel on H19-01 and are checking we see them here 31 Jan 2019
       
DAY 1 EMAIL from Mike - LED data and Es rsps

Hi Steph,
Guess what?
I have some more .fits data for thee!
Please seeth thine MOBY NAS at
/ftp1/Mike/L274/bs3cal
/ftp1/Mike/L274/bs4cal

Here is my first SNAFU:
could you please look at the one BS scan set from MOBY265 and verify
that the Es and LuBot tracks are reversed?
  Es was supposed to be on trk#2,
and LuBot was supposed to be on trk#13.  I think I probably mixed up the
jumper fibering inside the Splitter, such that Es = Trk#13, LuBot = trk#2.

It took me a couple days of fiddling to figure this out.  My troubles were:
1.) I don't have very refined image processing skills, and 2.) I had blue tape covering
the BS/RS Lu collector heads, which, is not "black" and leaks ambient light, and
3.) the lights in the tent are LEDs, with a spectral shape much like the LED Flashlight
I use to check MOBY sensors.  The good news is that I spent some time studying
Art's software's status-output for controlling shutters...

So, there are day01 data for both BS & RS that were LED Flashlight scans.
The day01 BS scans for LuBot were dark - because they were scanning the covered Es collector,
and the day01 RS scans for Es took 100 sec of exposure - because they were looking
at the blue-tape over the LuBot collector - with the tent lights on.
So, day01 BS Es & LuBot data were bogus.  All other day BS & RS data should be OK.

Day02 scans were BS & RS Es via cal lamp = FEL-F601(C3), but BS Es data were all dark scans,
because they were looking at the LuBot collector.  These had some signal when I had the
tent lights on during lamp warmup, but no signal when I darkend the tent to scan the FEL...
So day02 BS Es data were bogus.  Day02 RS Es data should be OK.

Day03 scans were BS Es & LuBot LED Flashlight scans = OK, and BS Es scans of F601(C3) = OK.

Luck!  MF

DAY 4 EMAIL from Mike - Lu rsps
Hi Steph,
Last night was pos-MOBY265 cal = BS & RS Lu via OL425(L10).
Zee data are at:
/ftp1/Mike/L274/bs3cal
/ftp1/Mike/L274/rs4cal

It will be challenging to tune the responses of the BS & RS
via iris'es and/or ND filters in the splitter, such that
exposure times are compatible for Es & Ed & Lu over 2x spectrographs!

For example, last night the exposure times (sec) :
LuBot,  BS=40,  RS=3
LuMid,  BS=43, RS=4.5
LuTop,  BS=16,  RS=16


So, last night I collected MANY more RS scan sets than BS scan sets for LuB & LuM,
but the same number of scan sets for BS & RS via LuTop.

I don't know that taking MANY more RS scan sets than BS scan sets was useful,
however, I thought I'd do it this time to maybe look at the variability of the
lamp via short-exp-time vs via long-exp-time - which might be more interesting
in the water via sunlight vs tent + lamp!

Q's:
FileZilla continues to challenge me.  It will not upload to the NAS for more than
aboot 20 or 30 files before crashing.  Then, sometimes, it will try to re-connect to the
NAS and re-start uploads - which is never successful.  So I have to Server/Disconnect
the File/Exit and re-start FileZilla & figure out where to re-start uploads.
This is the (repetative) error I get during file transfer failure:
Response:    150 Opening ASCII mode data connection for d20181204079.fits
Response:    425 Unable to build data connection: Operation not permitted
Error:    File transfer failed

Q1: Have you ever tried / do you have trouble FTPing large numbers of files
since you are now outside the MLML network/firewall kine?
Q2: Do you know how to clear old "Queued files" from an old session?

MF
DAY 5 EMAIL from Mike - Es wavecals
Hi Steph,
There are day5 data @ thine NAS for thee.
Day5 twas just the HgA & Ne lamps via Es collector.
I was not sure how many HgA lines would make it to the RS,
so I added the Ne lamp just to be safe.

The 1st Ne scan set was saturated @ the RS04, so I moved the lamp back
from the Es diffuser & repeated both BS & RS scan sets via Ne.

Note: BS03's Es is @ trk#13, and RS04's Es is @ trk#2

Tonight I plan to get all 5x wave lamps via LuTop = trk#6 on both BS & RS.
Then tonight/tomorrow the Hga & Ne via LuBot = trk#2 & 13.
Cheers, MF

DAY 6 EMAIL from Mike - LuTop, Bot wavecals
Hi Steph,
Day6 data are heading toward the NAS in starts & stops.
They were via 5x wave lamps @ LuTop = trk#6 for both BS3 & RS4,
plus HgA & Ne lamps @ LuBot = trk#2 for BS3  &  trk#13 for RS4.

Warning: all the Ne scans sets had saturated data...
Luck, MF


DAY 7 EMAIL from Mike - LuTop inttime cal
Hi Steph,
Well, the VPN connection is SLOW, but it was steady = no errors!
Thank you for your help setting this up.

Day07 was exposure time cals for the BS & RS.
The 2x filled-out log sheets for the BS apply to the RS (not filled out completely).
The OL455 photodiode Amp output was logged to file = 2018120701_OL455log.txt,
which is in the bs3cal dir, with the BS putty logs.

Let me know if this looks OK!  MF